# Nothing Conformal about Adaptive Conformal Inference

### v2

@@ 4 Nothing conformal is required
For all $t\in\mathbb{N}$, $\varepsilon_{t}\in[-\delta,1+\delta]$, almost surely, if ACI (15) is used together with a set predictor that outputs $\emptyset$ is predicted for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$.
@@
Assume that the sequence $\{\varepsilon_{t}\}_{t\in\mathbb{N}}$ is such that $\inf_{t}\varepsilon_{t}<-\gamma$ (the case when $\sup_{t}\varepsilon_{t}>1+\gamma$ is identical). By (15), $\sup_{t}|\varepsilon_{t+1}-\varepsilon_{t}|=\sup_{t}\gamma|\varepsilon-\text{err}_{t}^{\varepsilon_{t}}|<\gamma$. Thus, with positive probability, we may find $t\in\mathbb{N}$ such that $\varepsilon_{t}<0$ and $\varepsilon_{t+1}<\varepsilon_{t}$. However, by assumption, and (15),
$\varepsilon_{t}<0\implies\Gamma_{t}^{\varepsilon_{t}}=\boldsymbol{Y}\implies\text{err}_{t}^{\varepsilon_{t}}=0\implies\varepsilon_{t+1}=\varepsilon_{t}+\gamma(\varepsilon-\text{err}_{t}^{\varepsilon_{t}})\geq\varepsilon_{t}.$ (19)
@@
It is clear that no property of conformal, or even confidence predictors, is needed to achieve (16).
5 Why nested prediction sets matter
This section argues that, while not strictly necessary for achieving the finite sample guarantee (16) of ACI, we should restrict ourselves to confidence predictors, as this is the most general type of predictor that satisfy natural assumptions on prediction sets at a confidence level.
5.1 Validity is easy
We have seen that the same guarantee holds for ACI used with a predictor that outputs random subsets of $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. This predictor is not a confidence predictor because we have dropped the requirement on nested prediction sets. But if we are willing to dispense with nested prediction sets, we could do even better. The coin flip predictor outputs either $\emptyset$ or $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. Which one is output is determined by flipping a biased coin.
Definition 4 (Coin flip predictor).
A coin flip predictor is a set predictor that for each significance level $\varepsilon\in(0,1)$ outputs the prediction set
@@ 5.1 Validity is easy
The point is that the coin flip predictor is exactly valid. Its error sequence is a realisation of a sequence of independent Bernoulli variables, with parameter $\varepsilon$ for each $\varepsilon\in(0,1)$. However, it is clearly not very informative, and we should be hesitant to use it in practice.
We argue that if we want to predict with confidence, the very least we can ask for is a confidence predictor. Imagine that we play the following game: Before throwing and $n$ sided die, a player is allowed to bet that the outcome will be contained in a subset of ${1,\dots,n}$. Surely, we would not be willing to bet a larger amount on a smaller subset. Dropping the property of nested prediction sets is analogous to this game in the sense that the predictor could state a higher confidence in a smaller prediction set, which is absurd. If we are to mean anything reasonable by stating a confidence level, we simply must require that the prediction sets are nested. For this reason, we argue that, while ACI does not strictly require it, we should restrict its usage to extended confidence predictors in the sense of Definition 3 (and the analogous extended conformal predictors). But that raises a natural empirical question.
@@ 5.2 Do we gain anything from conformal predictors?
We have shown that ACI achieves the finite sample guarantee (16) for non-exchangeable data even for non confidence predictors, but we have argued that it would be absurd to drop the property of nested prediction sets. A natural question is whether using ACI with a conformal predictor has any advantages over using it together with a general confidence predictor. For non-exchangeable data, conformal predictors are not necessarily valid, which is also true for confidence predictors.
6 Experiments
We conduct three numerical experiments: one on synthetic data for a regression task, and two on real-world, non-exchangeable datasets, where we tackle both a regression and a classification task. These experiments aim to compare the performance of adaptive conformal inference (ACI) using conformal predictors versus confidence predictors under different data structures.
To evaluate the regression tasks, we employ the mean Winkler interval score (IS) [13], which assesses the quality of the prediction intervals. This score measures both the width of the interval and any penalties for miscoverage. The Winkler interval score for an individual prediction interval is calculated as follows. Let $l$ and $u$ be the lower and upper bounds of the prediction interval for $y$ at the desired significance level $\varepsilon$. The interval score for $[l,n]$ is
$S(l,u,y,\varepsilon):=\begin{cases}(u-l)+\frac{2}{\varepsilon}(l-y),&\text{if $y<l$}\\ u-l,&\text{if $y\in[l,u]$}\\ (u-l)+\frac{2}{\varepsilon}(y-u),&\text{if $y>u$}.\end{cases}$ (21)
The mean Winkler interval score is the mean value of the individual interval scores. Smaller values are better. Since there is a possibility of outputting infinite intervals, we compute interval scores only for finite prediction intervals, and report the fraction of infinite intervals separately. The Winkler interval score is a strictly proper scoring rule [14], meaning in essence that the lowest score is attained if the predicted intervals match the underlying distribution.
In the classification setting, we report the observed excess (OE) of the prediction sets [2], which is defined as
i.e. the average number of false labels included in the prediction sets at significance level $\varepsilon$. It is clear that smaller values are preferable. It is shown in [2] that the observed excess is a conditionally proper efficiency criterion, which in essence mean that the true conditional probability of the label, given the object, is always an optimal conformity score with respect to the conditionally proper efficiency criterion. In this sense, a conditional proper efficiency criterion is analogous to a proper scoring rule. A full discussion on efficiency criteria for conformal classifiers is beyond the scope of this work. For a detailed discussion, see Chapter 3 in [2].
In our evaluation, we have to modify the definition of OE slightly to account for the ACI update (15). We want to achieve the finite sample guarantee (16) with a target error rate $\varepsilon$. Then
$OE_{ACI}(\Gamma,\varepsilon):=\frac{1}{n}\sum_{i=1}^{n}|\Gamma^{\varepsilon_{i}}_{i}\backslash\{y_{i}\}.$ (23)
The difference is that we predict at significance level $\varepsilon_{i}$. This is a natural modification, and we will still refer to it as the observed excess (OE).
6.1 Synthetic data
Our first experiment, based on synthetic data from prior work [4], examines three distinct data settings:
@@ 6.1 Synthetic data
    Setting 1: i.i.d. data. Generate $N=2000$ i.i.d. examples $z_{i}=(x_{i},y_{1})$ with $X_{i}\sim\mathcal{N}(0,I_{4})$ and $Y_{i}\sim X_{i}^{T}\beta+\mathcal{N}(0,1)$ with coefficient vector $\beta=(2,1,0,0)$.
@@
    Setting 2: change points. Same setup as in Setting 1, but the coefficient vector changes. The first 500 points have coefficient vector $\beta_{1}=(2,1,0,0)$. The next 1000 points have coefficient vector $\beta_{2}=(0,-2,-1,0)$, and the last 500 points have coefficient vector $\beta_{3}=(0,0,2,1)$.
  • 3.
    Setting 3: distribution drift. Same setup but the coefficient vector changes from $\beta_{1}=(2,1,0,0)$ to $\beta_{n}(=0,0,2,1)$. The coefficient vector is computed by linear interpolation.
For each setting, we implement the following two methods with the desired error rate set to $\varepsilon=0.1$. For both methods, the significance level for the next prediction is computed using ACI. We ensure that both methods are extended confidence predictors in the sense of Definition 3 by simply outputting $\mathbb{R}$ for $\varepsilon\leq 0$ and $\emptyset$ for $\varepsilon\geq 1$.
@@
    Conformalized least squares (CP). We use the conformalized ridge regression algorithm [2] with ridge parameter set to 0 (which corresponds to least squares). The conformalized ridge regresison algorithm is described in A.1.
@@
    Prediction intervals from ordinary least squares (OLS). The ordinary least squares algorithm can output prediction intervals natively. We implement an online ordinary least squares algorithm, and output the least squares prediction intervals. The resulting confidence predictor is described in A.2.
For all three experiments on synthetic data, the first 100 examples are used as initial training set, and we run them 1000 times using different random seeds (ranging from 0 to 999). In all our experiments, we ask for a target error rate $\varepsilon=0.1$, and set $\varepsilon_{1}=\varepsilon$. Choosing the step size $\gamma$ according to (17) with $\delta=0.05$ would suggest $\gamma\approx 0.0096$.
Figure 1: Numerical simulation for synthetic data showing the average error rate and width of prediction intervals, averaged over 1000 independent trials. Gaps indicate infinite prediction intervals.
IID Changepoints Drift
Method Error Inf IS Error Inf IS Error Inf IS
CP 0.100 0 4.17 0.102 0.00660 9.41 0.104 0.000480 5.63
OLS 0.100 0 4.16 0.102 0.0300 9.28 0.104 0.00155 5.62
Table 1: Numerical simulation for synthetic data showing the average error rate, the fraction of infinite prediction intervals, and the average interval score for the finite intervals, averaged over 1000 independent trials. Inf denotes fraction of infinite prediction intervals, and IS is the mean Winkler interval score.
The results are shown in Figure 1 and Table 1. The first thing to note is that both methods achieve (16) as expected. The gaps in the right-hand side plots in 1 indicate an infinite prediction interval, and the fraction of such infinite intervals is presented in Table 1. We see that although the error rates are identical for both methods, CP outputs fewer infinite intervals in the second and third settings, with change points and drift. As can be seen in Table 1, the difference in terms of average interval score (IS) is not large, but CP outputs fewer infinite intervals on average.
The major difference is the computation time. On average, OLS was about eleven times faster on the same machine. As discussed in A, CP prediction sets can be computed in $\mathcal{O}(n\ln n)$, while $OLS$ require only $\mathcal{O}(n)$.
6.2 Real data
We perform two numerical experiments on publicly available real world datasets. First, a regression task where the aim is to predict the quality of wine, followed by a classification task of handwritten digits.
6.2.1 Wine quality
The Wine Quality dataset is publicly available through the UCI Machine Learning repository [15]. It consists of eleven features that may be useful for predicting the quality of the wine, which is encoded as an integer between 3 and 9, inclusive. Most labels are between 4 and 7. There are 6497 examples in total, 4898 white and 1599 red.
Exchangeability of the wine dataset has been studied in [16], where it is shown that red and white wines are not exchangeable. Thus, we use the white wines as training set, and predict the quality of the red wines.
Predicting the label can be seen either as regression or classification, and we choose to perform the regression task. For confidence predictor, we use the RandomForestQuantileRegressor from the quantile-forest python package [17]. We turn it into an inductive conformal predictor (ICP) by using the WrapRegressor class from the crepes python package [18]. We run our experiment 1000 times using different random seeds for each run. For the ICP, we must set aside a calibration set that is not used in training. Thus, we use 3/4 of the training set as proper training set, and the rest as calibration set. A different random split is used for each run. As a result, the training set for our confidence predictor (QRF) has size 4898 while the ICP has a proper training set size of just 3673 because the remaining data is used for calibration. Again, we choose step size according to (17) with $\delta=0.05$. With $N=1599$ we should choose $\gamma\approx 0.011$.
Figure 2: Numerical experiment on the Wine Quality dataset showing the average error rate and width of prediction intervals, averaged over 1000 independent trials.
Method Error IS Inf
ICP 0.103 5.05 0.00216
QRF 0.0925 4.41 0.0000382
Table 2: Numerical experiment on the Wine Quality dataset showing the mean error rate, interval scores for the finite prediction intervals, and the fraction of infinite intervals output. The results are averaged over 1000 independent trials. Inf denotes fraction of infinite prediction intervals, and IS is the mean Winkler interval score.
Our results are summarised in Figure 2 and Table 2. Again, both methods achieve the theoretical guarantee (16) as expected. Figure 2 summarises the mean running error rate, and the mean interval score (IS), averaged over the 1000 independent trials. It can be noted that CP is overly confident on average early on, while QRF is conservative. In terms of computational efficiency there is not much difference, ICP was introduced to overcome the often intractable computational cost of full CP, and the calibration step is quite fast. The mean error rate together with mean interval scores and the mean fraction of infinite intervals is presented in Table 2. It is worth pointing out that QRF produced infinite intervals at all in only eight out of 1000 trials, while CP did so in 696 trials.
6.2.2 USPS
The US Postal Service (USPS) dataset consists of 7291 training images and 2007 test images of handwritten digits from 0 to 9. The images are $16\times 16$ greyscale pixels. Exchangeability of the USPS dataset has been studied in [19, 20], and it is well known that the examples are not perfectly exchangeable. We use a RandomForestClassifier from the scikit-learn python package [21] with the default parameters, e.g. the forest consists of 100 trees. Again, we use the crepes python package [18], sacrificing 1/4 of the training data for calibration to train an ICP. We also define a confidence predictor by using the predict_proba method, which outputs estimated probabilities for every possible label (details in Appendix B). Choosing step size according to (17) with $\delta=0.05$ and $N=2007$ leads to $\gamma\approx 0.009$.
Figure 3: Numerical experiment on the USPS dataset showing the average error rate and size of prediction sets, averaged over 1000 independent trials.
Method Error Size OE
ICP 0.102 0.944 0.0457
RF 0.0848 0.100 0.0718
Table 3: Numerical experiment on the Wine Quality dataset showing the mean error rate, prediction set size, and observed excess (OE). The results are averaged over 1000 independent trials.
The results of the USPS experiment is summarised in Figure 3 and Table 3. As expected, both methods achieve the theoretical guarantee (16), but RF is quite a bit more conservative than CP, in particular early on, as can be seen in Figure 3. In terms of mean prediction set size, there is not much difference between the methods, but the OE of RF is relatively high, which is likely due to the large prediction sets output by RF in the beginning.
7 Conclusions
In this paper, we have demonstrated that adaptive conformal inference (ACI) does not rely on any specific properties of conformal predictors to achieve its finite sample guarantee (16). In fact, it does not even need the more general concept of a confidence predictor, where the prediction sets are required to be nested (see Definition 1). However, we have argued that the property of nested prediction sets is the very least one should require when predicting with a confidence level. Without it, the coin flip predictor (Definition 4) is exactly valid but rather unhelpful. Since the validity guarantees of conformal predictors are lost if the exchangeability assumption is violated, and ACI provides finite sample coverage guarantees of another kind, we asked if anything is gained by using ACI with a conformal predictor, over using it with a confidence predictor in situations when exchangeability can not be expected.
For our experiments with synthetic data, it seems that there is some advantage to using a conformal predictor insetting 2 and 3 with change points and drift, where the fraction of infinite intervals output by the extended confidence predictor based on ordinary least squares (OLS) was quite a bit larger that for conformalized least squares (CP). The main advantage of OLS lies in its less costly computation. It was about eleven times faster than CP. But these are online algorithms, where all available data is used for training.
When we use an inductive conformal predictor, some part of the training set has to be sacrificed for calibration. In the wine experiment, the confidence predictor RandomForestQuantileRegressor (QRF) performed better than the inductive conformal predictor (ICP) based on the same base algorithm, both in terms of mean interval score, and fraction of infinite intervals. Of 1000 independent trials, only eight contained infinite intervals, while ICP output infinite predictions in 696 trials. In the inductive setting, however, the computational disadvantage of CP disappears, as the calibration step required for ICP is quite fast. It does require sorting, but it only has to be performed once.
For the classification task in the USPS experiment, the picture is less clear. While the ICP based on random forest outperformed the confidence predictor based on the same algorithm in terms of observed excess, it is possible that better results could have been obtained by choosing a larger $\varepsilon_{1}$ to start the ACI iteration (15). It is clear from Figure 3 that the confidence predictor was way too conservative early on, and took some time to adapt. Much of the observed excess, i.e. the number of false labels included in the prediction sets, comes from the first 200 predictions, illustrating that there may be some utility in careful tuning of the initial control input $\varepsilon_{1}$.
We have seen that ACI with a confidence predictor can perform almost as well as with a conformal predictor in the online setting, while being significantly less costly in terms of computation. In the offline mode, a confidence prediction paired with ACI can sometimes outperform an ICP, which may be attributed to that the entire training set can be used to train, without sacrificing some training data for calibration. However, as we have seen, this is not always the case. If the training set is large enough, setting aside some training data for calibration may not be too detrimental. Our experiments are far from conclusive, and more research is needed to determine the relative advantages of ACI with conformal and confidence predictors. Future research should include finding good initial conditions for the control input, and experiments on more complex synthetic data with known properties. Our experiments were limited to using just one underlying machine learning method per dataset. It would be interesting to compare more methods to find out which methods are better suited as confidence predictors, than as ICPs. We speculate that it may come down to how much training data is required to achieve acceptable performance. If less is needed, there is no great loss in sacrificing some data for calibration.
Acknowledgement
The author acknowledge the Swedish Knowledge Foundation and industrial partners for financially supporting the research and education environment on Knowledge Intensive Product Realisation SPARK at Jönköping University, Sweden. Project: PREMACOP grant no. 20220187.
Appendix A Details on the online algorithms
@@
The least squares method goes back to Gauss and Legendre, and is a widely used regression method. Ridge regression is a generalisation of the basic procedure, dating back to the 1960s, introducing a non-negative ridge parameter $a$ (setting $a=0$ recovers ordinary least squares). In matrix form, it can be represented as
@@
where $Y_{n}:=(y_{1},\dots,y_{n})^{T}$, $X_{n}=(x_{1},\dots,x_{n})^{T}$, and the ridge regression prediction for an object $x$ is $\hat{y}:=x^{T}\omega$. Lest squares is the special case when $a=0$.
@@
A.1 Conformalised ridge regression
The conformalised ridge regression algorithm (CRR) is a combination of two algorithms; lower CRR and upper CRR, that produce the lower and upper bound respectively. For the upper CRR we use the nonconformity measure $y-\hat{y}$, and the lower uses $\hat{y}-y$, where $\hat{y}$ is the ridge regression prediction of $y$. The complete algorithm is presented in Algorithm 1.
@@ A.1 Conformalised ridge regression
In our experiments on synthetic data, we trivially extend Algorithm 1 by requiring the output to be (,)($\infty$,$\infty$)( - ∞ , ∞ ) for $\varepsilon\leq 0$ and $\emptyset$ for $\varepsilon\geq 1$, which results in an extended conformal predictor. It is shown in [22] that the prediction sets can be computed in $\mathcal{O}(n\ln n)$ in the online mode. Computing $A$ and $B$ can be done in time $\mathcal{O}(n)$, and sorting can be done in time $\mathcal{O}(n\ln n)$.
A.2 Ridge confidence predictor

### v4

@@ 5 The use of conformal predictors is not essential
For all $t\in\mathbb{N}$, $\varepsilon_{n}\in[-\delta,1+\delta]$, almost surely, if ACI (15) is used together with a set predictor that outputs $\emptyset$ for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$.
@@
Assume that the sequence $\{\varepsilon_{n}\}_{t\in\mathbb{N}}$ is such that $\inf_{n}\varepsilon_{n}<-\gamma$ (the case when $\sup_{n}\varepsilon_{n}>1+\gamma$ is identical). By (15), $\sup_{n}|\varepsilon_{n+1}-\varepsilon_{n}|=\sup_{n}\gamma|\varepsilon-\text{err}_{n}^{\varepsilon_{n}}|<\gamma$. Thus, with positive probability, we may find $t\in\mathbb{N}$ such that $\varepsilon_{n}<0$ and $\varepsilon_{n+1}<\varepsilon_{n}$. However, by assumption, and (15),
$\varepsilon_{n}<0\implies\Gamma_{n}^{\varepsilon_{n}}=\boldsymbol{Y}\implies\text{err}_{n}^{\varepsilon_{n}}=0\implies\varepsilon_{n+1}=\varepsilon_{n}+\gamma(\varepsilon-\text{err}_{n}^{\varepsilon_{n}})\geq\varepsilon_{n}.$ (19)
@@
It is clear that no property of conformal, or even confidence predictors, is required to achieve (16). As Lemma 1 and its proof demonstrate, the finite-sample guarantee (16) holds, provided that the predictor respects the boundaries ($\emptyset$ for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$), and the significance level is updated via (15). This guarantee arises solely from the behaviour of the iterative error control loop.
5.1 Why nested prediction sets matter
This section argues that while not strictly necessary for achieving the finite sample guarantee (16) of ACI, we should restrict ourselves to confidence predictors, as this is the most general type of predictor that satisfies natural desiderata for prediction sets at a confidence level.
5.1.1 Validity is easy without nested prediction sets
We have seen that the same guarantee holds for ACI used with a predictor that outputs random subsets of $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. This predictor is not a confidence predictor because we have dropped the requirement for nested prediction sets. However, if we are willing to dispense with nested prediction sets, we could do even better. The coin-flip predictor outputs either $\emptyset$ or $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. Which one is determined by flipping a biased coin.
Definition 4 (coin-flip predictor).
A coin-flip predictor is a set predictor that for each significance level $\varepsilon\in(0,1)$ outputs the prediction set
@@ 5.1.1 Validity is easy without nested prediction sets
The point is that the coin-flip predictor is exactly valid. Its error sequence is the realisation of a sequence of independent Bernoulli variables with parameter $\varepsilon$ for each $\varepsilon\in(0,1)$. However, this is clearly not very informative, and we should be hesitant to use it in practice.
5.1.2 Conflicting predictions
We argue that if we want to predict sets with confidence, the very least we can ask for is a confidence predictor whose prediction sets are nested, as illustrated in Figure 1. Imagine a model tasked with diagnosing patients based on symptoms, such as cough, shortness of breath, fever, and chest pain. The model outputs prediction sets at confidence levels $0.7,0.9$ and $0.99$, and the possible diagnoses are
  • 1.
    Healthy (no diagnosis)
  • 2.
    Pneumonia,
  • 3.
    Bronchitis,
  • 4.
    Asthma,
  • 5.
    Chronic Obstructive Pulmonary Disease (COPD).
Suppose that, for some patient, the model outputs the prediction sets in Table 1.
Confidence Prediction set
0.99 {Healthy, Pneumonia, Bronchitis, Asthma}
0.9 {Pneumonia, Bronchitis, Asthma}
0.7 {Healthy, COPD}
Table 1: A model that predicts with confidence, but does not have nested prediction sets.
The presented scenario illustrates a logical inconsistency in the model predictions. While asserting 90% confidence that the patient does not have a healthy diagnosis, the model simultaneously includes ”Healthy” within its 99% confidence prediction set. Furthermore, at the 70% confidence level, ”Healthy” reappears alongside ”COPD.” However, as the confidence level increases to 90% and subsequently 99%, ”COPD” is excluded. Given the potentially serious nature of these diagnoses, conflicting information from the model is problematic. Consequently, to ensure meaningful interpretation of confidence levels, the property of nested prediction sets is essential. Thus, we argue that if we are to mean anything reasonable by stating a confidence level, we simply must require that the prediction sets are nested. Therefore, while ACI does not strictly require it, its usage should be restricted to extended confidence predictors in the sense of Definition 3. However, this raises several empirical questions.
@@ 5.2 Do we gain anything from conformal predictors?
We have shown that ACI achieves the finite sample guarantee (16) for non-exchangeable data, even for non-confidence predictors, but we have argued that it would be absurd to drop the property of nested prediction sets. A natural question is whether using ACI with a CP has any advantages over using it together with an NCCP. For non-exchangeable data, CP is not necessarily valid, which is also true for NCCP.
The observation that ACI achieves the same finite sample guarantee for any confidence predictor, be it CP or NCCP, raises at least two questions regarding its application.
  1. 1.
    In the online setting, can we use ACI with an NCCP instead of a CP? The validity guarantee of CP is lost anyway, so the question is whether CP is more efficient (in the sense of outputting smaller prediciton sets) than NCCP.
  2. 2.
    In the offline setting, ICP requires sacrificing some training data for calibration. If the data are not exchangeable, does it even help to do so? Since the ACI finite sample coverage guarantee holds regardless, might it be better to use all available data for training and NCCP to be as efficient as possible?
We now proceed with our experiments, with the aim to answer these questions.
6 Experimental setup
Attempting to answer the questions raised by the observation that ACI achieves the same guarantee (16) for NCCP as for CP, we conduct four numerical experiments on publicly available datasets, known to be nonexchangeable.
@@ 6 Experimental setup
    The Wine Quality dataset is publicly available in the UCI Machine Learning repository [30]. It consists of 11 features that may be useful for predicting the quality of wine, which is encoded as an integer between 3 and 9, inclusive. Most labels are between 4 and 7. There are 6497 examples in total, 4898 white and 1599 red. The exchangeability of the wine dataset was studied by [31], where it was shown that red and white wines are not exchangeable.
@@
    The US Postal Service (USPS) dataset consists of 7291 training images and 2007 test images of handwritten digits from 0 to 9. The images are $16\times 16$ grayscale pixels. The exchangeability of the USPS dataset was studied in [32] and[33]. It is well known that the examples are not perfectly exchangeable.
6.1 Online experiments
Our online experiments attempt to answer the question of whether CP is more efficient than NCCP when used together with ACI in the online setting for nonexchangeable data. Recall that the ACI guarantee (16) holds for both CP and NCCP.
6.1.1 Online classification
We process the USPS dataset in the original order (which is known to be non-exchangeable) using the first 100 examples as the initial training set. The desired error rate is $\varepsilon=0.1$. We use the following two confidence predictors:
@@ 6.1.1 Online classification
    Nearest Neighbours Conformal Predictor (CP). The $k$-nearest neighbours CP is specified by the nonconformity measure
    $A(\sigma,(x,y))=\frac{\frac{1}{k}\sum_{x_{i}\in N_{k}(x,\sigma):y_{i}=y}d(x,x_{i})}{\frac{1}{k}\sum_{x_{j}\in N_{k}(x,\sigma):y_{j}\neq y}d(x,x_{j})}$ (21)
    where $\sigma=\Lbag(x_{1},y_{1}),\dots,(x_{n},y_{n})\Rbag\backslash\Lbag(x,y)\Rbag$, $d$ is the Euclidean distance, and $N_{k}(x,\sigma)$ denotes the set of the $k$ nearest neighbours of $x$ in $\sigma$ according to $d$ (or more generally, any distance).
@@
    Nearest Neighbours Confidence Predictor (NCCP). The $k$-nearest neighbours confidence predictor uses the confidence score $S_{k}(x,y)=N_{k}(x,y)/k$ to construct prediction sets. Here, $N_{k}(x,y)$ is the number of examples among the $k$ nearest neighbours of $x$ with label $y$. The prediction set is constructed using confidence thresholding: $\Gamma^{\varepsilon}(x)=\{y\in\boldsymbol{Y}:S_{k}(x,y)>\varepsilon\}$.
With 9198 test examples, and $\varepsilon_{1}=\varepsilon$ for simplicity, we aim for an absolute error deviation bound $\delta=0.01$, which, according to (17) means that we should choose the step size $\gamma\approx 0.001$.
The $k$-nearest neighbours predictor produces prediction sets by finding the $k$ nearest neighbours of the test point $x_{n}$ in the training data, which can be performed in time $\mathcal{O}(n)$. For the $k$-nearest neighbours conformal predictor, nonconformity scores must be computed for each example in the bag, which takes time $\mathcal{O}(n^{2})$, followed by computing the p-values for each potential label which takes time $\mathcal{O}(n)$ per label.
6.1.2 Online regression
Predicting the labels in the wine quality dataset can be seen as either regression or classification. We choose to perform the regression task again with the desired error rate $\varepsilon=0.1$ and using the first 100 examples as the initial training set. With 6397 test examples and $\varepsilon_{1}=\varepsilon$ for simplicity, we can achieve a guaranteed absolute coverage deviation $\delta=0.01$ if we choose $\gamma\approx 0.014$. We use the following confidence predictors:
  • 1.
    Conformalized Least Squares (CP). We use the conformalized ridge regression algorithm [3] with ridge parameter set to 0 (which corresponds to least squares). The conformalised ridge regression algorithm is described in Algorithm 1.
  • 2.
@@ 6.1.2 Online regression
    Least Squares Condfidence Predictor (NCCP). The ordinary least-squares algorithm can output prediction intervals natively. We implement an online ordinary least squares algorithm and output the least squares prediction intervals. The resulting confidence predictor is described in 2.
@@
The least-squares method goes back to Gauss and Legendre and is a widely used regression method. Ridge regression is a generalisation of the basic procedure, dating back to the 1960s, which introduces a non-negative ridge parameter $a$ (setting $a=0$ recovers ordinary least squares). In matrix form, it can be represented as
@@
where $Y_{n}:=(y_{1},\dots,y_{n})^{T}$ and $X_{n}=(x_{1},\dots,x_{n})^{T}$, and the ridge regression prediction for an object $x$ is $\hat{y}:=x^{T}\omega$.
@@
The conformalised ridge regression algorithm (CRR) is a combination of two algorithms: lower CRR and upper CRR, which produce lower and upper bounds, respectively. For the upper CRR, we use the nonconformity measure $y-\hat{y}$, and for the lower CRR, we use $\hat{y}-y$, where $\hat{y}$ is the ridge regression prediction of $y$. The complete algorithm is shown in Algorithm 1.
@@
In our experiments on synthetic data, we trivially extend Algorithm 1 by requiring the output to be (,)($\infty$,$\infty$)( - ∞ , ∞ ) for $\varepsilon\leq 0$ and $\emptyset$ for $\varepsilon\geq 1$, resulting in an extended conformal predictor. It was shown in [5] that the prediction sets can be computed in $\mathcal{O}(n\ln n)$ in the online mode. Computing $A$ and $B$ can be done in time $\mathcal{O}(n)$, and sorting can be done in time $\mathcal{O}(n\ln n)$.